# TrailBlazer: Trajectory Control for Diffusion-Based Video Generation

### 31 Dec 2023 v1

@@ 5 Conclusion
We have addressed the problem of controlling the motion of objects in a diffusion-based text-to-video model. Specifically, we introduced a combined spatial and temporal attention guidance algorithm operating in the pre-trained ZeroScope model. Our contributions are several. Firstly, the spatial location of a subject can be guided through simple bounding boxes. Secondly, bounding boxes and prompts can be animated via keyframes, enabling users to alter the trajectory and coarse behavior of the subject along the timeline. The resulting subject(s) fit seamlessly in the specified environment, providing a viable pipeline for video storytelling by casual users. Our approach requires no model finetuning, training, or online optimization, ensuring computational efficiency and a good user experience. Lastly, the results are natural, with desirable effects such as perspective, motion with the correct object orientation, and object/environment interactions arising automatically.
@@ 6 Implementation
In this section, we describe details of our implementation, including the core library, hyperparameters, and other pertinent information. Our method is developed using PyTorch 2.01 Paszke et al. (2019), and the Diffusers library version 0.21.4 from Huggingface Huggingface (2023). We override the the Diffusers pipeline TextToVideoSDPipeline to produce our implementation.
Parameters are selected as follows: We use classifier-free guidance with a strength of 9, conduct 40 denoising steps, and maintain a video resolution of 512x512 for the conventional stable diffusion backward denoising process. Regarding the parameters specific to our proposed method, the default values are as follows: We perform 5 editing steps for both spatial and temporal attention (denoted as $N_{S}\equiv N_{M}\equiv 5$). The editing coefficients $c_{m}\equiv 0.001$ and $c_{s}\equiv 0.1$ are used in both spatial and temporal attention in most cases. The number of trailing attention maps $|\mathcal{T}|$ is the only parameter that needs to be tuned. Generally, $10\leq|\mathcal{T}|\leq 20$ yields satisfactory results in practice.
As highlighted in section 1, we adapt the pre-trained ZeroScope [^4] Huggingface (2023):cerspense/zeroscope_v2_576w cerspense (2023) T2V model. This model is fine-tuned from the initial weights of ModelScope Luo et al. (2023) [^5] Huggingface (2023):damo-vilab/modelscope-damo-text-to-video-synthesis utilizing nearly ten thousand clips, each comprising 24 frames as training data. Consequently, we adhere to the recommended practice of setting the length of the synthesized sequence to 24 frames, drawing insights from user experiences shared in relevant blogs [^6] https://zeroscope.replicate.dev/.
@@ 7.2 Exploration and Ablation: Varied dynamic bbox sizes
Figure 13: Dynamic bbox sizes. The result showcases six synthesized video sequences with the subject directed by the yellow arrow starting at the position indicated by green bbox. The number of the bboxes (corresponding to the number of keyframes used in the experiment) is, clockwise from top-left, $|\mathcal{K}|=$ 3, 2, 2, 2, 2, and 2, respectively. The prompt used in each result: “The [X] swimming in the sea”, where “[X]” denotes the “fish” for the first and second rows, and “sardine” for the third row.
@@ 7.3 Exploration and Ablation: Speed control with multiple keys
Fig. 14 demonstrates controlling the subject’s speed through varying the number of keyframes in the video synthesis. Given the recommended sequence length $N_{f}=24$ for ZeroScope, we show the result of adding different keyframes in between the start and end keyframes at the left/right image boundary, simulating the cat running back and forth on the grass field. It is clear that the cat moves relatively naturally according to the motion flow indicated by the yellow arrows. For instance, the cat looks back first before turning around, rather than showing an unnatural motion where the position of the head and tail is instantaneously swapped. As the cat moves faster, motion blur also introduced in the result. We found that this motion blur is hard to eliminate using negative prompts.
@@ 7.5 Exploration and Ablation: Irregular trajectory
We illustrate irregular trajectories determined by varied keyframes in Fig. 16. The four experiments involve a zigzag trajectory (top-left), a triangle trajectory (top-right), a discontinuous trajectory (bottom-left), and a down-pointing triangle trajectory (bottom-right). In every result the horse shows high-speed running with motion blur. However, the results with turning points show limitations in depicting the horse quickly turning around and may show artifacts. For example, in the third frame of the down-pointing triangle case, the horse appears to swap its head and tail. Difficulty portraying this turn is somewhat expected, as horses cannot naturally execute tight high-speed turns, unlike cats or dogs. On the other hand, the down-pointing triangle video naturally introduces a perspective-like size change as the horse moves higher in the image, similar to the previous results in Fig. 13, and also the car/tiger example Fig. 6. In summary, maintaining consistency between the prompt and the timing and location of the keyframed bounding boxes is crucial for producing realistic results.
Figure 16: Irregular trajectory. The figure shows four synthesized video sequences with the horse subject directed according to the yellow arrows starting from the position indicated by green bbox. The frames highlighted in red indicates correspond to keyframes. The start and end keyframes are not indicated. The prompt used for all examples: “A horse galloping on the road”.

### 08 Apr 2024 v2

@@ 5 Conclusion
We have addressed the problem of controlling the motion of objects in a diffusion-based text-to-video model. Specifically, we introduced a combined spatial and temporal attention guidance algorithm, TrailBlazer, operating in the pre-trained ZeroScope model. The spatial location of a subject can be guided through simple bounding boxes. Bounding boxes and prompts can be animated via keyframes, enabling users to alter the trajectory and coarse behavior of the subject along the timeline. The resulting subject(s) fit seamlessly in the specified environment, providing a viable approach to video storytelling by casual users. Our approach requires no model finetuning, training, or online optimization, ensuring computational efficiency and a good user experience. Lastly, the results are natural, with desirable effects such as perspective, motion with the correct object orientation, and the interactions between object and environment arising automatically.
@@ 6 Implementation
In this section we describe details of implementation in TrailBlazer, including the core library, hyperparameters, and other pertinent information. Our method is developed using PyTorch 2.01 Paszke et al. (2019), and the Diffusers library version 0.21.4 from Huggingface Huggingface (2023). We override the Diffusers pipeline TextToVideoSDPipeline to produce our implementation.
Parameters are selected as follows: We use classifier-free guidance with a strength of 9, conduct 40 denoising steps, and maintain a video resolution of 512x512 for the conventional stable diffusion backward denoising process. For all the comparisons with Peekaboo Jain et al. (2023) we use their official repository [^3] https://github.com/microsoft/Peekaboo at the commit 6564274 (12 Feb 2024). In our comparisions we utilize a resolution of 576x320 as employed in the Peekaboo code to ensure fair assessment.
Regarding the parameters specific to our proposed method, the majority of our results are generated using the default values outlined as follows: We execute 5 editing steps for both spatial and temporal attention, denoted as $N_{S}\equiv N_{M}\equiv 5$. The editing coefficients $c_{w}\equiv 0.001$ and $c_{s}\equiv 0.1$ are used in both spatial and temporal attention in most cases. The number of trailing attention maps $|\mathcal{T}|$ is the only parameter that needs to be tuned. Generally, $10\leq|\mathcal{T}|\leq 20$ yields satisfactory results in practice and we set $|\mathcal{T}|\equiv 15$ for our paper results.
As highlighted in Sec. 1. in the main text, we adapt the pre-trained ZeroScope [^4] Huggingface (2023):cerspense/zeroscope_v2_576w cerspense (2023) T2V model. This model is fine-tuned from the initial weights of ModelScope Luo et al. (2023) [^5] Huggingface (2023):damo-vilab/modelscope-damo-text-to-video-synthesis utilizing nearly ten thousand clips, each comprising 24 frames as training data. Consequently, we adhere to the recommended practice of setting the length of the synthesized sequence to 24 frames, drawing insights from user experiences shared in relevant blogs.  [^6] https://zeroscope.replicate.dev/
@@
Following Jain et al. (2023), the following prompt set is used for the experiments in our quantitative comparison in Sec. 4.2 in the main text. We include it here only for completeness. The prompt word(s) in bold case is the subject for positioning:
  • A woodpecker climbing up a tree trunk.
  • A squirrel descending a tree after gathering nuts.
  • A bird diving towards the water to catch fish.
  • A frog leaping up to catch a fly.
  • A parrot flying upwards towards the treetops.
  • A squirrel jumping from one tree to another.
  • A rabbit burrowing downwards into its warren.
  • A satellite orbiting Earth in outer space.
  • A skateboarder performing tricks at a skate park.
  • A leaf falling gently from a tree.
  • A paper plane gliding in the air.
  • A bear climbing down a tree after spotting a threat.
  • A duck diving underwater in search of food.
  • A kangaroo hopping down a gentle slope.
  • An owl swooping down on its prey during the night.
  • A hot air balloon drifting across a clear sky.
  • A red double-decker bus moving through London streets.
  • A jet plane flying high in the sky.
  • A helicopter hovering above a cityscape.
  • A roller coaster looping in an amusement park.
  • A streetcar trundling down tracks in a historic district.
  • A rocket launching into space from a launchpad.
  • A deer standing in a snowy field.
  • A horse grazing in a meadow.
  • A fox sitting in a forest clearing.
  • A swan floating gracefully on a lake.
  • A panda munching bamboo in a bamboo forest.
  • A penguin standing on an iceberg.
  • A lion lying in the savanna grass.
  • An owl perched silently in a tree at night.
  • A dolphin just breaking the ocean surface.
  • A camel resting in a desert landscape.
  • A kangaroo standing in the Australian outback.
  • A colorful hot air balloon tethered to the ground.
7 Ablations
We conduct ablation experiments on the number of trailing attention maps and the number of temporal steps.
Trailing attention maps. Fig. 8 shows an ablation varying the number of trailing attention maps used in our spatial cross attention process, where the top row shows our method without trailing attention maps (e.g, $|\mathcal{T}|=0$) to the bottom row (e.g., $|\mathcal{T}|=30$). The guided bbox is annotated with green bbox moving from left to right. It is observed that the astronaut remains static at the image center without the trailing attention maps. In contrast, the synthesis with a large number of trailing attentions can lead to failed results such as a flag rather than the intended astronaut. A good number of edited trailing attention maps is between $|\mathcal{T}|=10$ and $|\mathcal{T}|=20$.
Temporal attention editing. We further show an ablation test in Fig. 9 with varied number of temporal attention editing steps. We take the case of the astronaut from Fig. 8 with $|\mathcal{T}|=10$, and set $N_{M}=0$ (no editing steps), and $N_{M}=10$. The result with $N_{M}=0$ shows a red blob moving from left to right. The value $N_{M}=10$ gives satisfactory result on the astronaut, but the background along the bbox path is missing. From these results we see that a reasonable balance between spatial and the temporal attention editing must be maintained, while extreme values of either produce poor results. An intermediate value such as $N_{M}=5$ used in most of our experiments produces the desired result of an astronaut moving over a moon background.
8 Limitations
Our method shares and inherits common failure cases of the underlying diffusion model. Notably, at the time of writing, models based on CLIP and Stable Diffusion sometimes generate deformed objects and struggle to generate multiple objects and correctly assign attributes (e.g. color) to objects. We show some failures in Fig. 10. For instance, we requested a red jeep driving on the road but the synthesis shows it sinking into a mud road. The panda example shows the camera moving instead of the panda itself. The red car has implausible deformation, and Darth Vader’s light saber turns into a surf board. The length of the resulting video clips is restricted to that produced by the pre-trained model, for instance, the 24 images in the case of ZeroScope. This is not a crucial limitation, as movies are commonly (with some exceptions!) composed of short “shots” of several seconds each. The bbox guides object placement without precisely constraining it. This is an advantage as well, however, since otherwise the user would have to specify the correct x-y aspect ratio for objects, a complicated task for non-artists.
Figure 10: Failure cases. Prompts used in subfigures: “A red jeep driving on the road”, “A red car driving on the highway”, “a panda eating bamboo”, and “Darth Vader surfing in waves”, where the bold prompt word is the directed subject.
9 Further comparison with Peekaboo
This section provides a comprehensive comparison between TrailBlazer and Peekaboo. In Fig. 11, we explore additional experiments encompassing various scenarios related to bbox size, location, and their combinations under extreme conditions. From top row to bottom: 1) Extremely fast motion by the timing of the second keyframe; 2) Rapid size changes along the bbox trajectory; 3) Zigzag trajectory; 4) Extremely fast motion through numerous keyframes; and 5) Extremely small bbox. Please refer to our the supplementary video to examine the motion for each of the following figures.
Figure 11: Extreme comparison: Various conditions. TrailBlazer (left) and Peekaboo (right) use identical bboxes, while T2V-Zero uses the corresponding motion vector instead. The same prompt is used across each method. The five prompts used from top: An elephant walking on the moon; a photorealistic whale jumping out of water while smoking a cigar; A horse galloping fast on a street; A dog is running on the grass; A clownfish swimming in a coral reef. The first column at left displays the bbox, and its trajectory. For the sequences with complex motion (2nd, 3rd, 4th row), the frames shown in the figure are denoted by the red dots along the trajectory in the first column. The orange bbox in the first row represents the starting motion of the elephant running to the right near the end of video clip. For additional details, please refer to the accompanying text.
As depicted in Fig. 11, TrailBlazer excels in most extreme scenarios for the synthesized subject’s location, motion speed, and identity. For example, our representation of an elephant maintains a stationary position for the initial 75% of the video before initiating movement then runs to the right. The whale gracefully descends into the ocean during the latter part of its jumping motion. The horse accurately follows a zigzag path, simulating a galloping motion. Remarkably, the dog seamlessly follows a large number of keyframes (8 keyframes) within a 24-frame video clip, covering the distance from one boundary to the opposite in approximately 2 frames. The clownfish fits into a tiny bbox. These successes are generally not evident in the Peekaboo Jain et al. (2023) method.
The metrics presented in Table. 3 are derived from the analysis of experiments shown in Fig. 11 using the AnimalKingdom dataset Ng et al. (2022), as described in our main text. The mIoU and FID in TrailBlazer surpasses Peekaboo, indicating that our method excels at effectively generating the subject in extreme conditions. Notably, as shown in Table. 3, our mIoU is approximately twice the value of Peekaboo. As mentioned in the main text, we believe the FVD in TrailBlazer is worse than Peekaboo in this section because the AnimalKingdom dataset does not contain the varied and extreme motion that we used in our experiments.
Method FID($\downarrow$) FVD($\downarrow$) IS($\uparrow$) KID($\downarrow$) mIoU($\uparrow$) CLIPSim($\uparrow$)
Peekaboo Jain et al. (2023) 258.34 2863.77 1.25 $\pm$ 0.24 6.71% 0.19 30.89
TrailBlazer 215.37 3416.41 1.47 $\pm$ 0.58 5.08% 0.39 29.41
Table 3: Quantitative results for static bbox.
10 Subject Morphing
Subject morphing involves blending semantics for generating images and videos. To our best knowledge, TrailBlazer is first in demonstrating subject morphing by prompt manipulation in the video diffusion domain. Related concepts have earlier been shown for image generation in MagicMix Liew et al. (2022) with, for example, the "corgi coffee machine".
While subject morphing through prompt embedding interpolation may seem less intuitive for real-world applications, it is widely used in the entertainment industry for example for superheroes (e.g., the She-Hulk can transform from a human to a monstrous character). For general usage, it could potentially serve as an entry point for generating new content that is more efficient than using a single prompt, particularly due to the limitations of CLIP Radford et al. (2021). For example, it might be challenging to generate a “fish-like” cat using the prompt "A fish-like cat walking on the grass" with a diffusion model. Instead, it would be easier to accomplish this goal by combining the prompt embeddings from “A fish swimming on the grass” and “A cat walking on the grass.”
Fig. 12 illustrates the morphing outcomes generated by TrailBlazer. All results are generated using default hyperparameter settings and involve linear interpolation of the prompt embeddings across video frames. The animated bounding boxes shift from right to left in the top two rows, and from left to right in the bottom two rows.
The outcome depicted in Fig. 12 demonstrates that morphing in a video clip can transition smoothly from one identity to another without significant artifacts. Notably, it avoids unrealistic deformations such as generating new joints in unexpected body parts (e.g., a tail on the head) or transforming one animal feature into another (e.g., an eye into an ear). Additionally, the subjects follow exactly the same motion in the synthesis (e.g., walking) across video frames.
Figure 12: Subject Morphing. The prompts used starting from the first row: “A [cat $\rightarrow$ dog] walking on the grass”, “A [cat walking $\rightarrow$ fish swimming] on the grass”, “A [parrot $\rightarrow$ king penguin] walking on the beach”, and “A [tiger $\rightarrow$ elephant] walking in the wild park.”. Please refer to the text for more detail.
@@ 11.2 Exploration and Ablation: Varied dynamic bbox sizes
Figure 15: Dynamic bbox sizes. The result showcases six synthesized video sequences with the subject directed by the yellow arrow starting at the position indicated by green bbox. The number of the bboxes corresponding to the number of keyframes used in the experiment is, clockwise from top-left, $|\mathcal{K}|=$ 3, 2, 2, 2, 2, and 2, respectively. The prompt used in each result: “The [X] swimming in the sea”, where “[X]” denotes the “fish” for the first and second rows, and “sardine” for the third row.
@@ 11.3 Exploration and Ablation: Speed control with multiple keys
Fig. 16 demonstrates controlling the subject’s speed through varying the number of keyframes in the video synthesis. Given the recommended sequence length $N_{f}=24$ for ZeroScope, we show the result of adding different keyframes in between the start and end keyframes at the left/right image boundary, simulating the cat running back and forth on the grass field. It is clear that the cat moves relatively naturally according to the motion flow indicated by the yellow arrows. For instance, the cat looks back first before turning around, rather than showing an unnatural motion where the position of the head and tail is instantaneously swapped. As the cat moves faster, motion blur also introduced in the result annotated with red arrows. We found that this motion blur is hard to eliminate using negative prompts.
@@ 11.5 Exploration and Ablation: Irregular trajectory
We illustrate irregular trajectories determined by varied keyframes in Fig. 18. The four experiments involve a zigzag trajectory (top-left), a triangle trajectory (top-right), a discontinuous trajectory (bottom-left), and a down-pointing triangle trajectory (bottom-right). In every result the horse shows high-speed running with motion blur. However, the results with turning points show limitations in depicting the horse quickly turning around and may show artifacts. For example, in the third frame of the down-pointing triangle case, the horse appears to swap its head and tail. Difficulty portraying this turn is somewhat expected, as horses cannot naturally execute tight high-speed turns, unlike cats or dogs. On the other hand, the down-pointing triangle video naturally introduces a perspective-like size change as the horse moves higher in the image, similar to the previous results in Fig. 15, and also the tiger example Fig. 6 in our main text. In summary, maintaining consistency between the prompt and the timing and location of the keyframed bounding boxes is crucial for producing realistic results.
Figure 18: Irregular trajectory. The figure shows four synthesized video sequences with the horse subject directed according to the yellow arrows starting from the position indicated by green bbox. The frames highlighted in red correspond to keyframes. The start and end keyframes are not indicated. The prompt used for all examples: “A horse galloping on the road”.